Help
3. A medication cup contains 45.0 milliliters (mL.) of a solution. How many tablespoons are there in the medication cup? 4. A tumor was found in the patients body with a diameter of 3 inches (in). What is the diameter of the tumor in millimeters (mm)? 5. A doctor orders an oral solution for your patient at 0.75 mg daily. You read the label on the bottle and find the dosage unit at I mL and the dose on hand at 1 mg. How may milliliters would you prepare for your patient? 6. The physician orders medication MS at 750 mg every 12 hours. You read the label on the bottle and find the dosage unit at 5 mL and the dose on hand at 125 mg. Determine the number of ml. you would give to the patient. 7. How many 200 mg tablets of medication TZ contain 1.6 g of medication TZ? 8. The order is for medication XX at a dose of 0.75 mg daily. The label on the bottle provides a dose on hand of 125 mcg (0.125 mg) per capsule. How much would you give the patient?
Answer #1

3.

1 Tablespoon = 15 ml

Formula: Milliliters(mL) = Tablespoon (tbsp) x 15

45ml = X (tbsp) x 15

X (tbsp) = 45ml/ 15ml = 3 tbsp

Hence, 45ml of solution contains 3 tablespoons

4.

1 inch is equal to 25.4 millimeters.

1" = 25.4 mm

then, 3" x 25.4 = 76.2 mm

Therefore, Tumor with 3 inches in diameter is equal to 76.2 millimeters

5.

Order: 0.75mg
quantity: 1ml
available dose: 1mg

Formula : Desireddose xQuantity Doseonhand

  

  \frac{0.75}{1} X 1= 0.75ml

The nurse will prepare 0.75 ml for the patient.   

6.

order: 750mg
available: 125mg
Quantity: 5ml

Formula : \frac{Desired dose}{Dose on hand} X Quantity

\frac{750}{125} X 5= 30ml

The nurse will give 30ml of the drug every 12th hourly.

7.

1 g = 1000 mg

Now convert 1.6 g to mg

1.6 x 1000 = 1600 mg

1.6 g = 1600mg

1600 mg / 200mg = 8

Therefore, 8 tablets of 200mg TZ contains 1.6g TZ

8.

order: 0.75mg

Available dose: 0.125mg

quantity :1

Formula : \frac{Desired dose}{Dose on hand} X Quantity

0.75 / 0.125 x1 = 6

6 capsules will be given to the patient.